Differentiate y= cos (6x +2)


1
Expert's answer
2022-03-21T03:56:40-0400

Let us differentiate "y= \cos (6x +2)." It follows that

"y'= -\sin (6x +2)\cdot (6x +2)'=-6\sin (6x +2)."